6. You receive a prescription written for a topical solution which calls for 45 g to be dispensed. The recipe for the solution is as follows: (How much of each ingredient is needed to produce 45 grams of topical solution?)
Betamethasone valerate 121 mg
Alcohol (95%) 20 ml
Propylene glycol 20 ml
Pluronic F127 22 g
Purified water qs 100 g
1
Expert's answer
2012-09-25T11:16:57-0400
Density of alcohol: 785.06 kg/m^3
Density of propylene glycol: 965.27
kg/m^3
The total weight in the recipe is:
Wr = 121 + 20 * 0.78506 + 20 *
0.96527 + 22000 + 100000 = 122126.1544761342
So, for the coefficient k =
45000 / 122126.1544761342 =
0.3684714399878518 having:
Betamethasone
valerate 121 *
0.3684714399878518 =
44.585044239 mg
Alcohol
(95%) 20
*
0.3684714399878518 = 7.3694288 ml
Propylene
glycol 20
*
0.3684714399878518 = 7.3694288 ml
Pluronic
F127 22
*
0.3684714399878518 = 8.10637168 g
Purified water
qs 100 *
0.3684714399878518 = 100 *
0.3684714399878518 g